2. Average Senior High School annual cost of tuition fee for all private schools last year


was Php43,700. A random sample of costs this year for 45 private schools indicated


that the sample mean was Php45,800 and a sample standard deviation was Php5,600.


At 0.01 level of significance, is there sufficient evidence to conclude that the cost


increased?

1
Expert's answer
2022-04-04T16:40:41-0400

"\mu =43700"

n=45

"X=45800"

"\sigma=5600"

"Z=\frac{X-\mu}{\frac{\sigma}{\sqrt{n}}}=\frac{45800-43700}{\frac{5600}{\sqrt{45}}}=2.52"

"Z_{\lambda/2}=Z_{0.01/2}=Z_{0.005}=2.807"

"2.52<2.807"

There is no sufficient evidence to conclude that the cost has increased.